Given that the y -intercept of a line is (0,-5) and the slope is 1/3, find the equation of a line in slope intercept form?
- What is the x- intercept?
- Find a point on the line.

Answer by richwmiller(9135)

y=x/3-5
if y=0 then x/3=5
x=15
(15,0) is the x intercept
the x and y intercepts are points on the line.
but here is a third point
y=3/3-5
y=1-5
(3,-4)